# Cybersecurity in Critical Infrastructures: A Post-Quantum Cryptography Perspective

### v1

@@ V Post-Quantum cryptography in Critical Infrastructure
Another important issue of integrating PQC in ICS/CI infrastructure is that it is also necessary to think about a resiliency solution, with the capability to adapt for different cryptography algorithms, since it is not exactly known if they would be secure against quantum attacks or specific cybersecurity requirements could be imposed by different countries and OT environments. Note, for example, the recent controversy regarding the calculations of the security level of Kyber-512 posed by Daniel Bernstein [143]. In principle, it seems rather difficult that such flexibility can be achieved by implementing those cryptographic protocols with a hardware solution. As an example, if the network of a electricity provider is secured by integrating a specific PQC protocol in hardware, were the case that such method is not reliable anymore, then all those chips introduced in the network should be substituted by new ones that implement the alternative. It is straightforward to see that such scenario would led to a high cost in terms of money and man power. Furthermore, as we exposed before, there are many PQc standardisation processes along the word. The generalized increment in the public investment in quantum technologies, Table 2, is a signal that countries consider the development of quantum technologies a priority. Moreover, this table refers only to public expend and it do now show private and/or military investment in quantum technology. It is a signal that in the near future each country could adopt their own PQC cryptosystem protocols and will require to fulfill their cybersecurity requirements to companies that operate/sell in their territory. Note that France has recently established a normative requiring cybersecurity for the communication protocols within critical infrastructures [38] and, even if no specifics on PQC are required yet, it is a matter of time that they do. Hence, having flexible PQC solutions would allow fulfilling the specific requirements imposed by those global agents.
@@
Regarding current PQC algorithms, those mainly focus on achieving quantum security to provide confidentiality, integrity and authenticity, being the latter two the most important characteristics; due to the fact that they are thought mainly from the point of view of IT. For those to bee implemented in OT, they should assure high availability and adaptability as discussed in section II.2. Several countries are trying to standardize PQC algorithms taking into account the perspective of IT systems.
@@
However, since ICS/CI have to be secure under quantum attacks as well, it is necessary to study how those algorithms work on OT networks and see which of them achieves all the requirements or if new PQC algorithms should be designed. Despite of the fact that there are many PQC families and many different protocols for each family, it is probable that none of them fulfill all the demanding requirements in CI environments. So, there could not be only a lack of PQC experimental work in OT environments, but also a lack of theoretical framework. If such were the case, cryptography we should urge crytpographers to find other possible families or protocols that fit those conditions.
VI Conclusion
@@
We have concluded that although there are many different PQC alternatives that seem to provide good security against quantum attacks for IT services in the near-future, their implementation in CI is not a trivial problem. The lack of security notions for the PQC families, the long lifetime span of OT devices, the fact that the communications within ICS have stringent requirements and that those are mainly composed of legacy elements of little computational capabilities imply that there is a current gap in terms of PQC protocols that can be seamlessly adapted to such scenarios. Moreover, the absence of a general benchmark of PQC algorithms under the same conditions (e.g. same processor for latency tests) makes it hard to make a top view comparison among them to conclude which could be well suited for implementation in CI networks. This is really important since cryptosystems that introduce too much latency reduce the availability of communication protocols, which could produce fatal consequences not only to the specific industry which suffers the shutdown, but also to all the interconnected industrial chain due to cascade effects. Therefore, we urge the post-quantum cryptography community to realize experimental studies which compare the existing protocols under similar conditions to find out suitable protocols and to think about possible PQC algorithms that may be tailored to fullfil the stringent conditions of OT services. Furthermore, we have discussed that PQC solutions for ICS/CI networks should be flexible in order to avoid huge economic and manpower costs if the implemented protocol results to be deprecated or if new governmental requirements are imposed since, for example, hardware solutions would imply the substitution of a humongous amounts of elements introduced in different points of a network that could be enormous in terms of space, i.e. in of the order of hundred of kilometers.
@@ Acknowledgements
Appendix A Tables
Lattice-based Public Key Private key Ct/Signature
NTRU [144, 101] 930 1234 930
NTRU-HRSS [145, 101] 1 138 1 450 1 138
NTRU-Prime [145, 146, 101] 994 15 158 897
NTRU+ [72, 147] 864 1 728 864
Falcon [148, 101] 897 7 553 666
FatSeal [74] 2 321 385 2 048
Peregrine [75, 147] 897 7 553 666
SOLMAE [76, 147] 897 7 553 666
CRYSTALS-Kyber [134, 101] 800 1 632 768
Saber [145, 101] 672 832 736
FrodoKEM [78, 101] 9 619 19 888 9 720
LAC PKE [79] 544 1 056 704
Aigis-Enc [80, 149] 672 1 568 672
AKCN-MLWE (AES256) [81] 10 560 10 560 8 610
TALE [82, 150] 736 1 504 704
AKCN-E8 (AES256) [83] 928 928 896
SCloud (AES256) [84] 98 800 19 800 82 300
TiGER [147, 86] 480 177 640
SMAUG [147, 85] 174 672 768
Crystal-Dilithium(AES192) [101] 1 312 2 528 2 420
Aigis-Sig [88] 672 1 568 672
Mulan (no access) [89]
GCKSign [90, 147] 1 760 288 1 952
NCC-Sign [91, 147] 1 440 2 400 2 529
HAETAE [92, 147] 992 1 376 1 463
Table 3: Performance of Lattice-based PQC protocols. Ciphertext (Ct) and keys length are expressed in Bytes. All values are taken from the level I security defined by NIST: ”Any attack that breaks the relevant security definition must require computational resources comparable to or greater than those required for key search on a block cipher with a 128-bit key (e.g. AES128)” [132]
Hash-based Public Key Private key Ct/signature
SPHINCS+ [62, 134, 101] 32 64 7 856
Table 4: Performance of Isogeny-based PQC protocols. Ciphertext (Ct) and keys length are expressed in Bytes. All values are taken from the level I security defined by NIST: ”Any attack that breaks the relevant security definition must require computational resources comparable to or greater than those required for key search on a block cipher with a 128-bit key (e.g. AES128)” [132]
Code-based Public Key Private key Ct/Signature
BIKE [101] 1 540 280 1 572
PALOMA [102, 147] 319 488 7 808 32
Cl. McEllice [101] 261 129 6 492 128
HQC [101] 2 249 40 4 481
Piglet-1 [135] 1 212 32 1 801
ROLLO-I [106, 147] 1 240 120 620
REDOG [107, 147] 14 250 1 450 830
Enhanced PQsigRm [108, 147] 2 000 000 1 032
Table 5: Performance of Code-based PQC protocols. Ciphertext (Ct) and keys length are expressed in Bytes. All values are taken from the level I security defined by NIST: ”Any attack that breaks the relevant security definition must require computational resources comparable to or greater than those required for key search on a block cipher with a 128-bit key (e.g. AES128)” [132]
Multivariate Public Key Private key Ct/Signature
GeMSS [134, 101] 352 168 16 33
Rainbow [112, 101] 157 800 611 300 164
MQ-Sign [114, 147] 328 441 15 561 134
Table 6: Performance of Multivariate PQC protocols. Ciphertext (Ct) and keys length are expressed in Bytes. All values are taken from the level I security defined by NIST: ”Any attack that breaks the relevant security definition must require computational resources comparable to or greater than those required for key search on a block cipher with a 128-bit key (e.g. AES128)” [132]
Isogeny-based Public Key Private key Ct/Signature
SIKE [101] 330 374 346
FIBS [124] 32 64 17 088
Table 7: Performance of Isogeny-based PQC protocols. Ciphertext (Ct) and keys length are expressed in Bytes. All values are taken from the level I security defined by NIST: ”Any attack that breaks the relevant security definition must require computational resources comparable to or greater than those required for key search on a block cipher with a 128-bit key (e.g. AES128)” [132]
Others Public Key Private key Ct/Signature
IPCC (AES80)[131] 4 800 400 92 000
AIMer [128] 32 16 5 904
Table 8: Performance of Isogeny-based PQC protocols. Ciphertext (Ct) and keys length are expressed in Bytes. All values are taken from the level I security defined by NIST: ”Any attack that breaks the relevant security definition must require computational resources comparable to or greater than those required for key search on a block cipher with a 128-bit key (e.g. AES128)” [132]

### v2

@@ V Post-Quantum cryptography in Critical Infrastructure
Another important issue for integrating PQC in ICS/CI infrastructure is that it is also necessary to think about a resiliency solution, with the capability to adapt for different cryptography algorithms, since it is not exactly known if they would be secure against quantum attacks  [^10] Recall the new proposal for solving LWE with quantum computers in polynomial time [185]. or specific cybersecurity requirements could be imposed by different countries and OT environments. Note also the recent controversy regarding the calculations of the security level of Kyber-512 posed by Daniel Bernstein [202] and the recently proposed method for breaking the Rainbow PQC cryptosystem in one weekend requiring a single laptop [148]. In principle, it seems rather difficult that such flexibility can be achieved by implementing those cryptographic protocols with a hardware solution, which is the most explored one for low latency solutions as shown in Section IV. As an example, if the network of a electricity provider is secured by integrating a specific PQC protocol in hardware, were the case that such method is not reliable anymore, then all those chips introduced in the network should be substituted by new ones that implement the alternative. It is straightforward to see that such scenario would lead to a high cost in terms of money and man power. Hybrid cryptography could be a good solution for maintaining classical cybersecurity even if the implemented PQC algorithm is proven to be insecure, but it will not be secure against quantum computers until the industry upgrade its PQC communication protocol. Furthermore, as we exposed before, there are many PQC standardisation processes along the word. The generalized increment in the public investment in quantum technologies, Table IX, is a signal that countries consider their development a priority. Moreover, this table refers only to public expend and it does not show private and/or military investment in quantum technology. It is a signal that in the near future each country could adopt their own PQC cryptosystem protocols and will require to fulfill their cybersecurity requirements to companies that operate/sell in their territory. Note that France has recently established a normative requiring cybersecurity for the communication protocols within critical infrastructures [50] and, even if no specifics on PQC are required yet, it is a matter of time that they will. Hence, having flexible PQC solutions would allow fulfilling the specific requirements imposed by these global agents.
Moreover, it is important to note that the PQC solutions integrated in this networks should be low-power and autonomous in terms of energy. This is related with the previous discussion on the required flexibility. Note that if the power of the PQC solutions require to be changed regularly (for example by using batteries), such requirement would also result in huge costs for the industry. Therefore, it is essential that the solutions are powered by in the same way as the other elements of the network. This is why they should be low-power as not to make the power system to be saturated, i.e. they should not be a problem for the power system of the network [203]. Also, some environments are more challenging than other (e.g. an oil extraction plant), so self powering methods may also be required as a function of the ICS to secure. For example, energy harvesting methods could be required to power some of the devices as in some IoT sensor networks [204], implying low-power consumption requirements.
@@
Regarding current PQC algorithms, those mainly focus on achieving quantum security to provide confidentiality, integrity and authenticity, being the first one the most important characteristic; due to the fact that they are thought mainly from the point of view of IT. For those to be implemented in OT, they should assure high availability and adaptability as discussed in Section II-B. Several countries are trying to standardize PQC algorithms taking only into account the perspective of IT systems, nonetheless, considering also the OT context is of pivotal importance for the security and safety of the industry.
@@
Subsequently, since ICS/CIs have to be secure under quantum attacks as well, it is necessary to study how those algorithms work on OT networks and see which of them achieve all the requirements. Despite of the fact that there are many PQC families and many different protocols for each family, it is probable that none of them fulfill all the demanding requirements in CI environments. So, there could not be only a lack of PQC experimental work in OT environments, but also a lack of theoretical framework. If such were the case, cryptography should urge cryptographers to find other possible families or protocols that fit those conditions before a future comes in which operational technologies have no protection to quantum attacks.
VI Conclusion and a guidelines
@@
Problems Importance Section
Latency OT protocols require low latency communications [47, 27, 192] II-A
Legacy devices PQC require a great amount of memory and computational resources [51, 53, 52] II-A
Data integrity and availability Data integrity and availability are the most important characteristics in OT cryptography II-A
Regulation The solution has to be resilient to fulfill the regulations of the different countries [48, 49, 50] IV-H
Agility PQC primitives security are not proved and HW solutions are not flexible IV
Theory There are not theoretical PQC specific proposals for OT protocols V
Experiments More tests of PQC algorithms in CI environments have to be done V
TABLE X: In this table we present the most valuable insights and problems for implementing PQC security in CI
@@
We have concluded that although there are many different PQC alternatives that seem to provide good security against quantum attacks for IT services in the near-future, their implementation in CI is not a trivial problem. The lack of security notions for the PQC families, the long lifetime span of OT devices, the fact that the communications within ICS have stringent requirements and that those are mainly composed of legacy elements of little computational capabilities imply that there is a current gap in terms of PQC protocols that can be seamlessly adapted to such scenarios. Moreover, the absence of a general benchmark of PQC algorithms under the same conditions (e.g. same processor for latency tests) makes it hard to make a top view comparison among them to conclude which could be well suited for implementation in CI networks. This is really important since cryptosystems that introduce too much latency reduce the availability of communication protocols, which could produce fatal consequences not only to the specific industry which suffers the shutdown, but also to all the interconnected industrial chain due to cascade effects. Therefore,we consider that the following points stand as some of the most relevant future research lines regarding this topic:
  • Conduct experimental studies comparing different PQC families under the same conditions. As explained through the article, the latency of PQC protocols is provided for specific scenarios and implementations (e.g. different processors), implying that a straight comparison by means of the literature data would not be accurate. This type of studies would clarify which protocols could be more suitable for integration in industrial networks as well as providing information to the community.
  • Optimized PQC implementations for OT networks should be investigated in order to understand the capabilities of current proposals. As stated before, there are not many fair comparisons of state-of-the-art PQC protocols to understand which could be potentially implementable in industrial networks. However, the optimization of those existing protocols to be dedicated to such scenarios should also be investigated. Those could be candidates for OT communication systems would there be any successful protocols to fill all the conditions required.
  • Propose PQC protocols from the point-of-view of the stringent conditions of OT sevices. As discussed, the problem of securing industrial networks is fundamentally different to the one of protecting IT systems. For example, in OT the aim should be authenticity/integrity rather than confidentiality. This with the very important requirement of low latency. In this sense, researcher on PQC proposals could think of their methods to target this problematic instead of protecting confidentiality, which is the usual target.
  • Propose flexible solutions for PQC integration in ICS/CI networks. In Section V, we discussed that PQC solutions for ICS/CI networks should be flexible in order to avoid huge economic and manpower costs if the implemented protocol results to be deprecated or if new governmental requirements are imposed since, for example, hardware solutions would imply the substitution of a humongous amounts of elements introduced in different points of a network that could be enormous in terms of space, i.e. in of the order of hundred of kilometers. Thus, proposing, for example, programmable methods that are flexible enough to change protocols if required is important for this post-quantum transition.
  • Standardization processes for PQC implementations in industrial environments should be conducted. Through this document, many PQC standardization efforts around the world have been discussed, but those are oriented towards IT communications. As commented before, both communication scenarios are very different in the requirements for cryptography, indicating that the IT and OT implementations will diverge. Therefore, efforts regarding PQC implementations for OT networks should be pushed worldwide. Importantly, this should be done with a fast pace since, as discussed before, equipment that is not quantum secure could be vulnerable through their lifetime.